OpenStudy (anonymous):

Find the volume of a rectangular solid with the following dimensions: 2√3 cm x 7√5 cm x 4√6 cm.

OpenStudy (dumbcow):

volume = length*width*height just multiply the dimensions

OpenStudy (anonymous):

= 2 * 7 * 4 *√3 * √5 * √6 = 56 * √90 or 56* 3 √10 = 168 √10

OpenStudy (anonymous):

am getting \[(24\sqrt{2})(7\sqrt5)\]